OpenStudy (anonymous):

Alg. 1 multiple choice! helppp. What is the value of the range of the function f(x) = x² + 2 for the domain value 1/4? y = 2 y = 2 1/4 y = 2 1/8 y = 2 1/16

OpenStudy (anonymous):

\[f \left( \frac{ 1 }{ 4 } \right)=\left( \frac{ 1 }{ 4 } \right)^2+2\]

OpenStudy (anonymous):

what does f = ?

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

okay, so its 2 1/16?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thanks:) and can you hep me with the vetica line test thing?

OpenStudy (anonymous):

help*

OpenStudy (anonymous):

Use the vertical line test to determine whether the relation graphed below is a function. yes no cannot be determined

OpenStudy (anonymous):

nvm i got it. its yes.

OpenStudy (anonymous):

cannot be determined

OpenStudy (anonymous):

unless the points are connected

OpenStudy (anonymous):

@lyzzy7 do you get it?